OpenStudy (anonymous):

Find the derivative of the function h(x) =∫(upper=√x) (lower=1) s^2/(s^2+1)ds

hartnn (hartnn):

derivative of integration: \(\Large \dfrac{d}{dx}\int \limits_{f(x)}^{g(x)}h(x)dx = h(g(x))g'(x)-h(f(x))f'(x)\) use this!

hartnn (hartnn):

in your case, lower limit f(x) is a constant, so there won't be a 2nd term

hartnn (hartnn):

\(\Large \dfrac{d}{dx}\int \limits_{a}^{g(x)}h(x)dx = h(g(x))g'(x)\)

hartnn (hartnn):

need help using this ?

OpenStudy (anonymous):

1/(2x+2)

hartnn (hartnn):

how?

hartnn (hartnn):